1 Empire Ltd. needs Rs 1,000,000 to build a new factory which will yield EBIT of Rs .150,000 per year. The company has to choose between two alternative financing plans: 75 per cent equity and 25 per cent debt or 50 per cent equity and 50 per cent debt. Under the first plan shares can be sold for Rs 40 per share and the interest rate on debt will be 16 per cent. Determine the EPS 


1
Expert's answer
2021-02-10T15:04:02-0500

plan 1

Equity=750000

Debt=250000

Number of shares=750000/40

=18750

Interest rate=40000

=750000-250000-40000

Eps=460000/18750

=24.5

Plan 2

Equity=500000

Debt=500000

Number of shares=50000

Eps=75000/50000

=15